OpenStudy (anonymous):

For the function f(x) = -15x^4 +12, what is the value of f(x) when x=-2? F. -228 G. -108 H. 132 J. 252

OpenStudy (solomonzelman):

\(\large\color{blue}{ \bf f(x) = -15x^4 +12 }\) just giving couple examples. \(\large\color{blue}{ \bf f(3) = -15(3)^4 +12=-15(81)+12=-1215+12=1203 }\) \(\large\color{blue}{ \bf f(-1) = -15(-1)^4 +12=-15(1)+12=15+12=27 }\) \(\large\color{blue}{ \bf f(-2) = ~~you~~~~tell~~~~me. }\)

OpenStudy (anonymous):

I know how to do it it's just that I don't have a calculator can you please give me the answer

OpenStudy (solomonzelman):

\[(-2)^4=16,~~~~~~~~~~~~(-15)(16)=-240.\]

OpenStudy (anonymous):

-240x=-24